# WritingBench: A Comprehensive Benchmark for Generative Writing

### v3

@@ 3.2 Query-Dependent Evaluation Framework
Suri 4.97 4.4 5.5 5.6 5.3 5.0 4.1 5.0 5.1 4.8 5.2 5.0 5.4 4.5 4.0
LongWriter 7.91 7.9 7.9 8.0 8.1 8.1 7.7 8.1 7.6 7.9 8.2 8.1 8.8 7.7 7.7
Qwen-2.5-7B-filtered 8.49 8.6 8.4 8.4 8.4 8.6 8.4 8.6 8.5 8.6 8.8 8.5 9.0 8.5 8.6
Llama-3.1-8B-filtered 8.49 8.6 8.4 8.5 8.4 8.6 8.4 8.6 8.5 8.6 8.8 8.5 8.9 8.5 8.5
Table 3: WritingBench performance of different LLMs across 6 domains and 3 writing requirements evaluated with our critic model (scale: 1-10). The six domains include: (D1) Academic & Engineering, (D2) Finance & Business, (D3) Politics & Law, (D4) Literature & Art, (D5) Education, and (D6) Advertising & Marketing. The three writing requirements assessed are: (R1) Style, (R2) Format, and (R3) Length. Here, “C" indicates category-specific scores.
3.3 Critic Model
To alleviate the computational overhead with LLM-based evaluation, we develop a dedicated critic model, $\mathcal{M}$, designed to implement our rubric-based scoring framework. Specifically, this model performs the mapping $\mathcal{M}_{c}:(q,r,C_{i})\mapsto[1,10]\times\mathcal{J}$, where the output consists of a numerical score and corresponding justification text, $\mathcal{J}$, both in accordance with the predefined evaluation rubric.
We fine-tune the critic model on a dataset comprising 50K instances, which are collected using LLMs in our experiments. The dataset encompasses diverse queries, evaluation criteria, and model responses to enhance the robustness of evaluation. The experiments presented in Section 4.3 validate the consistency of the critic model.
@@ 3.4 Evaluation-Guided Data Curation for Writing Enhancement
The query-dependent evaluation framework enables systematic data curation across diverse writing tasks through two-phase filtering. First, applying instance-specific criteria generation to initial 24K SFT samples. Subsequently, the critic model filters 50% samples, yielding 12K framework-curated samples with higher quality. Fine-tuning experiments are conducted using the llama-3.1-8b-instruct and qwen-2.5-7b-instruct models. Evaluations are performed on two writing benchmarks to validate the framework’s effectiveness in identifying high-quality writing samples.
@@ 4 Experiment
Figure 5: Scores of different models across different subdomains in WritingBench. Red indicates higher score and blue refers to lower score. The figures are the average score of each subdomain for different models.
@@ 4.1 Experiment Settings
This section outlines the settings employed in our experiments using the WritingBench framework. The evaluation protocol, and the model training configurations are as follows:
Evaluation Protocol: Uniform generation settings are applied to model responses on WritingBench, allowing up to 16,000 tokens. We set the temperature at 0.7, with a top-k of 20 and top-p of 0.8, ensuring a balance of creativity. The critic model uses the same parameters, with input length capped at 2,048 tokens for scoring stability.
Model Training: The critic model is fine-tuned on the Qwen-2.5-7B-Instruct base model, using the AdamW optimizer with a 7e-6 learning rate. It is trained on 50K SFT data for 3 epochs across 8xA100 GPUs (batch size 64, 8-step accumulation). The writing models, trained on both Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct for 5 epochs on 32xA100 GPUs, achieving a total batch size of 128 with 4-step gradient accumulation.
@@ 4.2 Comparison between LLMs
We evaluate 16 LLMs: ChatGPT-4o-latest Hurst et al. (2024), o1-Preview Jaech et al. (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Pro Reid et al. (2024), Qwen-Max Team (2025), Deepseek-R1 DeepSeek-AI et al. (2025), Deepseek-V3 DeepSeek-AI et al. (2024), Mistral-Large-Instruct MistralAI (2024), Qwen-2.5-72B-Instruct and Qwen-2.5-7B-Instruct Yang et al. (2024), Llama-3.3-70B-Instruct and Llama-3.1-8B-Instruct Dubey et al. (2024), Suri Pham et al. (2024), LongWriter Bai et al. (2024), and our writing model, Qwen-2.5-7B fine-tuned model and Llama-3.1-8B fine-tuned model, on WritingBench. Each query is assessed by 5 instance-specific criteria using a 10-point scale, with subdomain-specific subcategory heatmaps revealing detailed variations.
Key Insights from Domain Scores: From Table 3, we observe that Education (D5) always yields high performance among models, followed by Academic & Engineering (D1). Literature & Art (D5) is the lowest-scoring domain, showing notable variance. CoT-capable models like Deepseek-R1 and o1-preview outperform their non-CoT counterparts models, indicating CoT’s potential in handling narrative and creative content. Figure 5 identifies challenges in niche areas such as novel continuations, tender proposals, and white papers where models generally score lower. These tasks demand a higher level of knowledge, long-text generation capabilities, and adherence to contextual consistency, pinpointing areas for further enhancement.
Key Insights from Requirement Scores: Table 3 presents scores for three requirement subsets: R columns showing overall scores and C columns showing requirement-specific averages. For example, if a style query includes two relevant criteria (as confirmed by human annotation), the C column averages scores for those criteria only. Style (R1) and Format (R2) dimensions reveal model distinctions , with Deepseek-R1 and Qwen-Max often excels others. Nevertheless, length requirements (C of R3) remains challenging, especially in section-specific constraints (see Appendix 16 for examples) and extended text production (see Section 4.5).
The overall analysis of WritingBench experiment highlights: Deepseek-R1 consistently leads across both domain and requirement dimensions, showcasing its versatility and strong language capabilities. There is significant performance variance within creative content domains, with CoT models showing notable promise—a hypothesis further confirmed by our comparison (see Appendix C.1, where the model incorporating CoT mechanisms outperform the one without CoT on creative writing tasks). Additionally, we observe that advanced models often achieve higher specialized scores (C column) for the three common requirements than their overall scores (R column). The criteria outside these specialized sets tend to emphasize content aspects, such as integration with source materials and writing depth, highlighting the need to improve content quality. This analysis benchmarks existing model capabilities and underscores specific areas needing improvement for future development.
@@ 4.3 Human Consistency
To validate the alignment between automated evaluation and human judgment, we conducted human evaluation on 300 queries, covering all 100 subdomains. Five professionally trained annotators with linguistic backgrounds perform pairwise comparisons of model responses. For each query, two responses are randomly selected from different models. Annotators select the preferred response or declared equivalence based on the query’s requirements. We compare our dynamic query-dependent criteria against two baselines: static globally uniform criteria and static domain-specific customized criteria (static criteria are designed by domain experts) on two LLM judges: ChatGPT-4o-latest and Claude-3.5-Sonnet.
Evaluation Metric Judge Score
Static Global ChatGPT 69%
Static Domain-Specific ChatGPT 40%
Dynamic Query-Dependent ChatGPT 79%
Static Global Claude 65%
Static Domain-Specific Claude 59%
Dynamic Query-Dependent Claude 87%
Dynamic Query-Dependent Critic Model 83%
Table 4: Comparison of human consistency scores across different criteria generation methods. ChatGPT corresponds to ChatGPT-4o-latest, Claude corresponds to Claude-3.5-Sonnet.
As shown in Table 4, our dynamic query-dependent criteria achieve superior human alignment compared to static, both globally uniform or domain-specific customized criteria. Notably, domain-specific criteria underperform despite customization, since our queries involve highly diverse tasks and varied sources. These findings confirm that context-sensitive query-dependent evaluation better captures real-world writing complexity compared to conventional static approaches. Furthermore, the critic model attains 83% agreement, confirming its practical viability.
@@ 4.4 Ablation of Data Curation for Writing-Enhanced Models
Leveraging the query-dependent evaluation framework described in Section 3.2, we perform instance-specific criteria generation an initial 24K SFT dataset spanning various domains, including long-input queries enriched with supplemental materials. Our critic model then apply a query-dependent evaluation to filter this dataset down to a high-quality 12K subset. We conduct fine-tuning experiments using the Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models. As shown in 5, both models, trained on the filtered data, demonstrate significant performance improvements over their previous versions. Notably, they outperform models trained on the full dataset and even approached the capabilities of SOTA models. Further validation on an alternative benchmark, LongBench-Write (following the settings of quality evaluation in Bai et al. (2024)), confirm these findings, with our models surpassing baseline performance.
These results validate the robustness of our query-dependent evaluation strategy and the utility of our critic model for curating high-quality writing samples. This integrated approach enables smaller models to compete with, and even surpass, larger models across diverse writing tasks.
Models WritingBench Benchmark2
Deepseek-R1 8.55 4.79
Qwen-2.5-7B-Instruct 7.43 4.39
Llama-3.1-8B-Instruct 6.35 3.12
Qwen-2.5-7B-all 8.46 4.69
Qwen-2.5-7B-filtered 8.49 4.70
Llama-3.1-8B-all 8.45 4.65
Llama-3.1-8B-filtered 8.49 4.65
Table 5: Performance evaluation of our writing models on two benchmarks.’-filtered’ indicates models trained with filtered data, while ’-all’ refers to those trained with the full dataset.
4.5 Ablation of Length
We assess the performance of LLMs across varying input and output lengths, with statistical validity ensured by excluding intervals containing fewer than 5 samples.As demonstrated in Appendix 6, experiments reveal that most SOTA models generally maintain consistent performance regardless of input length variations, attributable to their advanced long-context comprehension capabilities.
However, analysis in Appendix 7 shows that most models exhibit inherent limitations in response generation length, typically producing outputs constrained to approximately 3,000 tokens. While output quality remains relatively stable within this range, smaller models suffer more performance degradation characterized by repetitive outputs. Notably, only two models, Qwen-Max and LongWriter, which incorporate architectural optimizations for long-form generation, effectively support extended response lengths. It underscores the importance of enhancing long-output capabilities and length optimization in writing.
@@ 5 Conclusion
In this paper, we introduce WritingBench, a comprehensive benchmark designed to evaluate LLMs in generative writing across diverse domains. It includes 1,239 queries spanning 6 primary domains and 100 subdomains, providing evaluation dimensions for style, format, and length requirements. Our query-dependent evaluation framework, supported by a critic model, achieves high human alignment. Evaluation efficiency is further demonstrated by compact models trained on curated data, achieving performance comparable to SOTA models. By making WritingBench and its resources publicly available, we aim to foster further research and advancements in LLM writing capabilities.

### v4

@@ 3.2 Query-Dependent Evaluation Framework
Suri 3.20±0.042 2.5 3.8 3.6 3.5 3.0 2.5 3.2 3.6 3.2 3.7 3.1 3.2 3.0 3.0
LongWriter 6.27±0.081 6.2 6.4 6.4 6.4 6.3 6.0 6.5 6.0 6.3 7.4 6.3 6.7 6.3 6.8
Qwen-2.5-7B-filtered 7.44±0.058 7.7 7.2 7.4 7.2 7.5 7.3 7.7 7.7 7.5 8.4 7.6 8.1 7.4 7.2
Llama-3.1-8B-filtered 7.39±0.045 7.5 7.3 7.4 7.2 7.3 7.3 7.5 7.8 7.4 8.3 7.5 8.0 7.4 7.1
We identify three commonly encountered writing requirements: style, format, and length from real-world writing scenarios (frequency analysis detailed in Appendix A.1). For each requirement, we create two evaluation subsets. Two annotators review each query to classify each criteria in $C_{q}$ into style, format, length, or none. Based on their annotations, we perform final verification and define the subsets as follows: (1) The first requirement-involved subset includes all queries involving a specific requirement and their entire set of criteria. For example, if any criterion in $C_{q}$ relates to style, the query and its entire criteria set $C_{q}$ belong to this subset. Scores for all five criteria are calculated, corresponding to the R1/R2/R3 columns in Table 3. (2) The second category-specific subset includes only the criteria related to the specific requirement. For instance, if $c_{1}$ and $c_{4}$ in $C_{q}$ are format-related, then the query and the subset $c_{1},c_{4}$} belong to this subset. Only the scores for $c_{1},c_{4}$} are calculated, corresponding to the C columns. The first subset serves as an overall evaluation of writing quality for a specific requirement, while the second subset provides more targeted insights into performance on that particular capability.
Phase 2 - Rubric-based Scoring: For each criterion $c_{i}\in C_{q}$, the evaluator independently assigns a score on a 10-point scale to a response $r$, along with a justification. The final overall score is computed by averaging scores across all criteria. Scoring prompt is provided in Appendix C.6.
To alleviate the computational overhead with LLM-based evaluation, we develop a dedicated critic model, $\mathcal{M}$, designed to implement our rubric-based scoring framework. Specifically, this model performs the mapping $\mathcal{M}_{c}:(q,r,c_{i})\mapsto[1,10]\times\mathcal{J}$, where the output consists of a numerical score and corresponding justification text, $\mathcal{J}$, both in accordance with the predefined evaluation rubric. The critic model is fine-tuned on a dataset of 155K instances scored by Claude-3.7, which demonstrates higher human alignment consistency (as discussed in Section 4.3). When building the training data of scoring prompts, the queries and criteria are drawn from WritingBench. The response portion is generated using approximately 40 different models, including those evaluated in Section 4.2, as well as additional models of varying types and sizes, such as Claude-3.5-Haiku claude-3.5 and Llama-3.2-1B-Instruct llama-3.2. Finally, after removing instances where scoring failed, approximately 155K samples are constructed for training. The experiments described in Section 4.3 confirm the consistency of the critic model.
@@ 3.3 Evaluation-Guided Data Curation for Writing Enhancement
The query-dependent evaluation framework enables systematic data curation across diverse writing tasks through two-phase filtering. To validate its effectiveness in data curation, we conduct SFT experiments using constructed datasets. We utilize the secondary domain taxonomy of WritingBench and follow the first two steps outlined in Section 3.1.1 (removing the material generation suggestion part from the prompt in Appendix C.4, requiring only the refined query from the model) to let LLMs generate writing queries. To enrich query diversity, we employ two models, GPT-4o and Claude-3.5-Sonnet, to each generate 60 queries in both Chinese and English for every subdomain, totaling 24K queries. Responses are uniformly generated using an advanced model, Deepseek-R1.
Subsequently, we apply the query-dependent evaluation metric with our critic model described in Section 3.2 to select top-50% samples per subdomain, resulting in 12K high-quality data samples. Fine-tuning experiments are conducted using the Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models. Both models (Qwen-2.5-7B-filtered and Llama-3.1-8B-filtered) demonstrated significant performance improvements over their base versions and even outperformed larger models such as Llama-3.3-70B-Instruct and Qwen-2.5-72B-Instruct in our experiments on two benchmarks, confirming the framework’s effectiveness in identifying high-quality writing samples.
@@ 4 Experiment
Figure 5: Score heatmap of models across 100 subdomains in WritingBench. The figure shows the average score per subdomain for each model. Red indicates higher score and blue indicates lower.
@@ 4.1 Experiment Settings
This section outlines the settings employed in our experiments using the WritingBench framework. The evaluation protocol and the model training configurations are as follows:
Evaluation Protocol: Uniform generation settings are applied to model responses on WritingBench, with maximum sequence lengths capped at 16,000 tokens (or platform maximum if lower). We configure the generation process with a temperature of 0.7, a top-k value of 20, and top-p of 0.8, ensuring a balance of creativity and coherence. For both the critic model and LLM scoring, we use consistent configurations. The temperature is set to 1.0, in line with the Arena-Hard-Auto framework arena-hard. We apply a top-p value of 0.95 without any top-k filtering, and the maximum length is set at 2,048 tokens to ensure efficient and reliable output generation.
Model Training: The critic model is fine-tuned on the Qwen-2.5-7B-Instruct base model, using the AdamW optimizer with a 7e-6 learning rate. It is trained on 155K SFT data for 3 epochs across 16xA100 GPUs (batch size 64, 4-step accumulation). The writing models, trained on both Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct for 5 epochs on 32xA100 GPUs, achieving a total batch size of 128 with 4-step gradient accumulation.
Table 4: Comparison of human consistency scores across different criteria generation methods. Claude corresponds to Claude-3.7. Critic refers to the critic model.
Evaluation Metric Judge Score
Static Global GPT-4o 69%
Static Domain-Specific GPT-4o 40%
Dynamic Query-Dependent GPT-4o 79%
Static Global Claude 67%
Static Domain-Specific Claude 58%
Dynamic Query-Dependent Claude 87%
Dynamic Query-Dependent Critic 84%
Table 5: Performance of our writing models on two benchmarks. ‘-filtered’ indicates models trained with filtered data, while ‘-all’ uses the full dataset. Eval2 refers to LongBench-Write longwriter.
Models WritingBench Eval2
Deepseek-R1 7.70 4.79
Qwen-2.5-7B-Instruct 5.64 4.39
Llama-3.1-8B-Instruct 4.42 3.12
Qwen-2.5-7B-all 7.36 4.69
Qwen-2.5-7B-filtered 7.44 4.70
Llama-3.1-8B-all 7.34 4.65
Llama-3.1-8B-filtered 7.39 4.65
@@ 4.2 Comparison between LLMs
We evaluate 17 LLMs: GPT-4o [^1] In this paper, we specifically use GPT-4o, version gpt-4o-2025-01-29. (4ol), o1-Preview (o1), Claude-3.7 and its thinking version Claude-3.7-thinking (claude), Gemini-1.5-Pro (gemini), Qwen-Max (qwen-max), Deepseek-R1 (r1), Deepseek-V3 (deepseek-v3), Mistral-Large-Instruct (mistral), Qwen-2.5-72B-Instruct and Qwen-2.5-7B-Instruct (qwen2.5), Llama-3.3-70B-Instruct and Llama-3.1-8B-Instruct (llama), Suri (suri), LongWriter (longwriter), and our writing model, Qwen-2.5-7B-filtered (fine-tuned on Qwen-2.5-7B-Instruct) and Llama-3.1-8B-filtered (fine-tuned on Llama-3.1-8B-Instruct), on WritingBench. For models capable of reasoning, the reasoning content is excluded from the evaluated response. Each query is assessed by 5 instance-specific criteria using a 10-point scale scoring by our critic model. The overall scores of the models are presented in Table 3, along with their performance across two languages and various domain and requirement subsets. Detailed variations are further revealed through subdomain-specific subcategory heatmap in Figure 5.
Key Insights from Domain Scores: The Education (D5) and Advertising & Marketing (D6) domains always yield substantial performance across models. In contrast, the Academic & Engineering (D1) and Finance & Business (D2) domains present greater challenges, as these tasks inherently require more sophisticated information processing and integration capabilities. Our evaluation across 100 subdomains further identifies persistent difficulties in niche areas like writing bid proposals, financial reports, and white papers, where models generally achieve lower scores. These tasks demand a higher level of knowledge, long-text generation capabilities, and adherence to contextual consistency, pinpointing areas for further enhancement.
The Literature & Art (D4) domain exhibits notable performance variance among models. Reasoning-capable architectures such as Claude-3.7-thinking, Deepseek-R1, and o1-Preview outperform their non-reasoning counterparts, indicating the potential of CoT techniques in processing narrative and creative content. To further validate CoT’s efficacy in creative writing, we employ the 12k SFT dataset described in Section 3.3 to fine-tune Qwen-32b-Instruct using both CoT-integrated and non-CoT approaches. Evaluations are conducted on the Literature & Art (D4) subset of WritingBench and EQ-Bench eqbench, a specialized benchmark designed for creative writing tasks. The results reveal that the reasoning model consistently surpass both the baseline Qwen-32b-Instruct and its non-reasoning variants, with detailed results provided in Appendix A.2.
Key Insights from Requirement Scores: Most models perform well in the style dimension, followed by format, with length being the weakest. We observe that advanced models often achieve higher specialized scores (C column) for the three common requirements compared to their overall scores (R column). Criteria outside these specialized sets tend to emphasize content-related aspects, such as integration with source materials and writing depth, underscoring the need to improve content quality.
Length requirements remain particularly difficult, especially in section-specific constraints (see Appendix 7 for examples) and extended text generation. Additionally, we evaluate the performance of LLMs across varying input and output lengths (see Appendix A.4 for details). Advanced models generally sustain consistent performance across varying input lengths, leveraging their strong long-context comprehension abilities. However, regarding output length, most models demonstrate inherent limitations, typically capping their responses at around 3,000 tokens. In contrast, Claude-3.7 and its reasoning-enhanced version, as well as Qwen-Max, stand out for their capacity to generate extended outputs effectively. These results emphasize the critical need to improve long-output generation and refine length optimization in writing-related tasks.
The overall analysis of WritingBench experiment highlights: (1) Claude-3.7-thinking consistently leads across both domain and requirement dimensions, followed by its non-reasoning variant, showcasing versatility and strong language capabilities; (2) A significant performance variance is observed within creative content domains, where models incorporating CoT mechanisms surpass those without, demonstrating CoT’s potential in LLM writing; (3) Cross-lingual inconsistencies in models like Deepseek-R1 and Llama-3.3-70B-Instruct suggest limitations in multilingual knowledge alignment. Furthermore, detailed analysis on subdomains reveals persistent difficulties in knowledge-intensive tasks. Ablation experiments on output length emphasize that generating long-form outputs remains an obstacle for current models. This analysis not only benchmarks the existing capabilities of these models but also underscores specific areas needing improvement for future development. For the most up-to-date results, please refer to the online leaderboard [^2] For clearer display, online leaderboard scores use the same calculations but are rescaled to a 100-point scale..
@@ 4.3 Human Consistency
To validate the alignment between automated evaluation and human judgment, we conduct a human evaluation study involving 300 queries (constructed in the same pipeline as described in Section 3.1 but not included in WritingBench). Five professionally trained annotators with linguistic backgrounds perform pairwise comparisons of model responses. For each query, two responses are generated by two randomly selected models (drawn from the same set of 17 models used for evaluation in Section 4.2). The triplet <Query, Response_A, Response_B> is presented on the same page to enable direct comparison. Annotators are instructed to carefully read the query and select their preferred response or declare equivalence (A/B/Tie), with no criteria information provided to ensure unbiased decision-making (detailed instructions and the annotation interface are provided in Appendix A.3).
We then compare these human preferences with the scores assigned by our critic model to the responses — if the critic model rate Response_A higher and the annotators also prefer Response_A, it is counted as alignment. To further evaluate the effectiveness of our dynamic query-dependent criteria, we compare it against two baselines: static globally uniform criteria and static domain-specific customized criteria (designed by domain experts). This comparison was conducted using two LLM-based judges, GPT-4o and Claude-3.7. Both models score the responses using the rubric-based scoring method described in Section 3.2, using scoring prompts provided in Appendix C.6.
As shown in Table 4.1, our dynamic query-dependent criteria achieve superior human alignment compared to static, both globally uniform or domain-specific customized criteria. Notably, domain-specific criteria underperform despite customization, since our queries involve highly diverse tasks and varied sources. These findings confirm that context-sensitive query-dependent evaluation better captures real-world writing complexity compared to conventional static approaches. Furthermore, the critic model attains 84% agreement, confirming its practical viability.
@@ 4.4 Ablation of Data Curation for Writing-Enhanced Models
To validate the data curation capabilities of the query-dependent evaluation framework described in Section 3.2, we conduct fine-tuning experiments on two datasets: the initial 24K dataset constructed in Section 3.3 and the 12K subset curated using the query-dependent evaluation framework. We experiment with two models of different architectures, Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct, and evaluate their performance on two benchmarks: WritingBench and LongBench-Write, a general-purpose writing benchmark (following the quality evaluation settings outlined in (longwriter)).
As shown in Table 4.1, both models trained on the filtered 12K dataset demonstrate significant performance improvements over their previous versions. Notably, they outperform models trained on the full 24K dataset and even approach the capabilities of advanced models. These results validate the robustness of our query-dependent evaluation strategy and highlight the effectiveness of our critic model in curating high-quality writing samples. This integrated approach enables smaller models to compete with, and in some cases surpass, larger models across a wide range of writing tasks.
@@ 5 Conclusion
In this paper, we introduce WritingBench, a comprehensive benchmark designed to evaluate LLMs in generative writing across diverse domains. It includes 1,000 queries spanning 6 primary domains and 100 subdomains, providing evaluation dimensions for style, format, and length requirements. Our query-dependent evaluation framework, supported by a critic model, achieves high human alignment. Evaluation efficiency is further demonstrated by compact models trained on curated data, outperforming GPT-4o in writing. By making WritingBench and its resources publicly available, we aim to foster further research and advancements in LLM writing capabilities.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The methodologies for the contributions mentioned in the abstract and introduction are detailed in the subsections of Section 3, with experimental validation provided in Section 4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.